The balance sheet for Seuss Company at the end of the current fiscal year indicated the following:

Bonds payable, 10% (20-year term) $5,000,000
Preferred 10% stock, $100 par 1,000,000
Common stock, $10 par 2,000,000

Income before income tax was $1,500,000 and income taxes were $200,000 for the current year. Cash dividends paid on common stock during the current year totaled $150,000. The common stock sells for $75 per share at the end of the year.

Required:

Determine each of the following:

Round ratios and percentages to one decimal place, and monetary amounts to the nearest cent.

1. Times interest earned times
2. Earnings per share on common stock $
3. Price-earnings ratio
4. Dividends per share of common stock $
5. Dividend yield %
Answer #1

1) Interest expense = 5000000*10% = 500000

Times interest earned = Income before interest and tax/Interest expense = (1500000+500000)/500000 = 4 Times

2) Earning per share of Common Stock = (Income after tax-Income tax-preferred dividend)/Share outstanding = (1500000-200000-100000)/200000 = 6 per share

3) Price earning ratio = 75/6 = 12.50 times

4) Dividend per share of Common Stock = 150000/200000 = 0.75 per share

5) Dividend yield = 0.75/75 = 1%